OpenStudy (anonymous):

PLEASE HELP Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

@kelliegirl33 PLEASE HELLP MEEE

OpenStudy (tkhunny):

Why don't we start from the top and have you demonstrate the point-slope form. Go!

OpenStudy (anonymous):

I am so confused and I have to submit this assignment in like 20 minutes or I fail please just help me

OpenStudy (tkhunny):

I would love to help you. I need you to do the work. Point-Slope Form \((y-y_{0})=m(x-x_{0})\) Substitute the known point \((x_{0},y_{0}) = (3,-2)\) and the known slope (m = 2), and you'll be done with part A

OpenStudy (anonymous):

ok so (y- -2) = 2(x- 3)?

OpenStudy (tkhunny):

Done. Off to Part B. Just solve for y.

OpenStudy (anonymous):

(y- -2) = 2(x- 3) y- -2 = 2x- 6 +2 +2 y= 2x -4?

OpenStudy (tkhunny):

y - (-2) = y+2 Give it another go.

OpenStudy (anonymous):

im really confused

OpenStudy (anonymous):

I don't know what to do

OpenStudy (tkhunny):

Why. You just missed a piece. Go back and get it. (y- -2) = 2(x- 3) (y+2) = 2(x- 3) y + 2 = 2x - 6 There. You did everything correctly except that -(-2) = +2 Okay, now solve for y.

OpenStudy (anonymous):

how do I love for y

OpenStudy (tkhunny):

Just like you did before, but with the right numbers in there.

OpenStudy (anonymous):

y + 2 = 2x - 6 - 2 -2 ------------ y = 2x -4

OpenStudy (tkhunny):

-6 - 2 = -8

OpenStudy (anonymous):

ok so it y = 2x -8?

OpenStudy (tkhunny):

Perfect. Off to Part C What is Standard Form? ax + by = c or ax + by - c = 0

OpenStudy (anonymous):

the frist one

OpenStudy (anonymous):

ax + by = c

OpenStudy (tkhunny):

Okay, put it in that form. 1) Subtract y 2) Add 8 That will be it.

OpenStudy (anonymous):

2x = 8?

OpenStudy (tkhunny):

Where did the y go? y = 2x - 8 Subtract y 0 = 2x - 8 - y It doesn't disappear.

OpenStudy (anonymous):

so is that iin standard form

OpenStudy (tkhunny):

Add 8

OpenStudy (anonymous):

0 = 2x + 8 ?

OpenStudy (tkhunny):

Again, where did the y go? Why does it keep disappearing? y = 2x - 8 Subtract y 0 = 2x - 8 - y Add 8 8 = 2x - y Done!

OpenStudy (anonymous):

so im done?!

OpenStudy (tkhunny):

It's sad that you don't know. You might need to review this course. If you pass, please spend some time in review before challenging the next course. You will never have any fun if you are always struggling to catch up.

OpenStudy (anonymous):

bye